6x+5y=30
x=y+5
how do we solve and prove?

Answer by fractalier(6550)

From
6x + 5y = 30 and
x = y + 5
we should substitute the second equation into the first and get
6(y + 5) + 5y = 30
6y + 30 + 5y = 30
11y + 30 = 30
11y = 0 and
y = 0
x is the 5.